# Towards Generalized Video Quality Assessment: A Weak-to-Strong Learning Paradigm

### v3

@@ Abstract
Video quality assessment (VQA) seeks to predict the perceptual quality of a video in alignment with human visual perception, serving as a fundamental tool for quantifying quality degradation across video processing workflows. The dominant VQA paradigm relies on supervised training with human-labeled datasets, which, despite substantial progress, still suffers from poor generalization to unseen video content. Moreover, its reliance on human annotations—which are labor-intensive and costly—makes it difficult to scale datasets for improving model generalization. In this work, we explore weak-to-strong (W2S) learning as a new paradigm for advancing VQA without reliance on large-scale human-labeled datasets. We first provide empirical evidence that a straightforward W2S strategy allows a strong student model to not only match its weak teacher on in-domain benchmarks but also surpass it on out-of-distribution (OOD) benchmarks, revealing a distinct weak-to-strong effect in VQA. Building on this insight, we propose a novel framework that enhances W2S learning from two aspects: (1) integrating homogeneous and heterogeneous supervision signals from diverse VQA teachers—including off-the-shelf VQA models and synthetic distortion simulators—via a learn-to-rank formulation, and (2) iterative W2S training, where each strong student is recycled as the teacher in subsequent cycles, progressively focusing on challenging cases. Extensive experiments show that our method achieves state-of-the-art results across both in-domain and OOD benchmarks, with especially strong gains in OOD scenarios. Our findings highlight W2S learning as a principled route to break annotation barriers and achieve scalable generalization in VQA, with implications extending to broader alignment and evaluation tasks.
@@ 1 Introduction
Video quality assessment (VQA) [^1] This work focuses on no-reference (NR) or blind VQA, which assesses video quality without relying on additional reference information. (Min et al., 2024) plays an important role in modern video processing systems, delivering objective quality measurements used to optimize end-user Quality of
@@
Experience (QoE). With the advances in deep neural networks (DNNs) (He et al., 2016; Dosovitskiy et al., 2020; Liu et al., 2021) and the increasing availability of human-annotated datasets (Hosu et al., 2017; Sinno & Bovik, 2018; Wang et al., 2019; Ying et al., 2021), current VQA models (Wu et al., 2022; 2023a; 2023b; Sun et al., 2024) have achieved significant progress through supervised learning. Nevertheless, supervised learning inherently faces a limitation: the generalization of the VQA models heavily depends on the diversity of the training data. For example, even top-tier VQA models (Sun et al., 2024; Wu et al., 2022; 2023a; 2023b) exhibit significant performance drops in out-of-distribution evaluations, as illustrated in Fig. 1.
Figure 1: Significant performance drop of state-of-the-art models on out-of-distribution datasets.
@@
Existing VQA research has primarily focused on constructing scene-specific datasets (Li et al., 2019b; Madhusudana et al., 2021; Yu et al., 2022; Shang et al., 2023) or large-scale datasets (Götz-Hahn et al., 2021; Jia et al., 2025) to improve model generalization across different video content and distortions. However, constructing such datasets is highly resource-intensive. A standardized subjective experiment comprises two key phases: test sample curation and subjective quality annotation. The test sample curation phase necessitates rigorous selection of representative video samples, as inadequate sampling strategies risk producing oversimplified datasets (i.e.,easy dataset” problem (Sun et al., 2024; Cao et al., 2024)) and may induce model overfitting. Meanwhile, subjective annotation—though vital—is laborious and costly. International Telecommunication Union (ITU) standards (ITU-T P.910, 2008) outline specific recommendations for experimental setups, including display conditions, stimulus duration, subject count, and rating methodologies. These constraints, though necessary for statistically meaningful annotations, impede larger-scale dataset expansion due to prohibitive annotation costs.
@@
Therefore, these limitations naturally raise an important question: Can we train stronger VQA models without relying on large-scale human-annotated datasets? Prior efforts have investigated self-supervised and unsupervised VQA approaches (Chen et al., 2021b; a; 2022; Madhusudana et al., 2023; Mitra & Soundararajan, 2024) which primarily employ contrastive learning with proxy tasks such as distortion-type or severity classification on synthetically generated data. However, such methods face two key shortcomings: (1) they fail to capture high-level visual content and aesthetic characteristics that are critical for perceptual quality assessment, and (2) they inadequately model authentic distortion patterns in real-world videos, which often arise from complex nonlinear degradation processes. As a result, their performance still lags significantly behind supervised counterparts.
Recent progress in weak-to-strong (W2S) generalization provides a promising approach for tackling this open problem. In this paradigm, a strong student model—equipped with higher learning capacity or powerful pre-trained knowledge—can learn effectively from the supervision of a weaker model and further generalize to hard examples beyond the teacher’s reach. It is thus natural to leverage an existing VQA model as a weak teacher to distill a stronger one, obviating the need for human-annotated labels. This approach raises two critical questions: (1) How effectively does W2S generalization apply to VQA, a task that inherently involves subjective human perception rather than deterministic high-level semantics, and (2) How can we enhance its performance to meet the demands of practical VQA applications?
This work investigates these two problems. First, we empirically demonstrate that a straightforward W2S generalization approach enables the student model to match the performance of its weak teacher (e.g.,, off-the-shelf VQA models) on in-domain benchmarks and surpass it on out-of-domain (OOD) benchmarks, revealing a clear weak-to-strong generalization effect in VQA.
Second, we advance W2S learning for VQA from two aspects: integrating diverse supervision signals and iterative W2S training. For the former, we incorporate multiple types of “VQA models” as weak models to refine and diversify the supervised signals, including (1) ensembling homogeneous VQA models (i.e., off-the-shelf VQA models) to improve the reliability of supervision, and (2) integrating heterogeneous teachers (i.e., synthetic distortion simulators) to enrich the supervision space. To unify these heterogeneous supervision signals, we reformulate quality regression as a ranking problem to make the model to learn quality assessment capabilities through pairwise comparisons. For the latter, we propose an iterative W2S learning strategy with difficulty-guided sampling, where each trained strong model is recycled as the weak teacher for the next iteration. Within each cycle, we deliberately select difficult samples so that subsequent models focus on challenging cases beyond the reach of weaker teachers, thereby progressively expanding the generalization capacity of the student model.
@@
  • We empirically validate a distinct W2S generalization effect in VQA, providing a new paradigm for advancing self-supervised and weakly supervised approaches for VQA.
  • We introduce a novel W2S generalization framework that integrates heterogeneous supervision signals from diverse teachers and incorporates an iterative W2S training strategy.
  • Within this framework, our student model achieves state-of-the-art results on both in-domain and OOD benchmarks, with particularly notable gains on OOD performance.
@@ 2.1 VQA Models
Supervised VQA. Early VQA models (Saad et al., 2014; Mittal et al., 2015) were largely knowledge-driven, extracting handcrafted features (e.g., natural scene statistics (Mittal et al., 2012), motion cues (Konrad & Dubois, 1992)) to quantify distortions and training shallow regressors for quality prediction. Subsequent approaches (Li et al., 2019a; Ying et al., 2021) shifted to representation learning, employing pre-trained DNNs to extract frame-level quality representations, coupled with sequence models such as GRUs or Transformers for temporal regression. More recent efforts adopt end-to-end fine-tuning of advanced vision architectures, including Vision Transformers (ViTs) (Dosovitskiy et al., 2020) and large multimodal models (LMMs) (Wu et al., 2023b), with the designs such as grid-based mini-patch sampling or key-frame selection to mitigate the computational burden of full-video training. While these advancements have significantly improved the performance of VQA models on in-domain datasets, they still struggle to generalize satisfactorily to OOD datasets.
Self-supervised VQA. These methods primarily learn quality-aware representations through contrastive learning frameworks with proxy tasks such as next-frame feature discrimination and distortion type/severity classification (Chen et al., 2021a; Madhusudana et al., 2023), or via encoder–decoder reconstruction of pristine videos from distorted inputs (Xie et al., 2024). These representations are typically adapted for quality prediction by fine-tuning a lightweight linear projector with human-annotated labels. More recently, researchers (Wu et al., 2023a; Agnolucci et al., 2024) have explored leveraging the perceptual capability of vision–language models for zero-shot video quality assessment, for example by estimating the relative likelihood of prompts such as “high quality” versus “low quality.”
VQA as Ranking. Ranking-based methods reformulate quality prediction from a regression problem into a ranking problem. To this end, various loss functions such as hinge loss (Liu et al., 2017), fidelity loss (Zhang et al., 2021), binary cross-entropy loss (Zhu et al., 2024), and differentiable approximations of Spearman Rank Correlation loss (Li et al., 2022) have been employed to learn relative quality rankings from pairwise comparisons or groups of samples. Such methods are particularly effective in mitigating the misalignment of quality scales across different datasets and can be applied in scenarios where only relative quality labels are available. Consequently, they have been widely adopted in weakly supervised training and mixed-dataset training. In this work, we also adopt a learning-to-rank strategy to unify the heterogeneous supervisory signals provided by diverse weak teachers.
2.2 Weak-to-strong Generalization
Weak-to-strong (W2S) generalization studies how strong models can learn from weaker supervision yet surpass their teachers. Early empirical studies (Burns et al., 2023) showed that simply fine-tuning a strong model on weak labels already allows the student to outperform its weak teacher across domains such as NLP, reward modeling, and games. Building on these foundations, subsequent studies have focused on improving the quality of weak supervision. Co-supervised and mixture-of-experts approaches (Liu & Alahi, 2024) combine diverse weak teachers to mitigate noise and bias; ensemble and scalable oversight methods (Sang et al., 2024) enhance teacher reliability through aggregation and debate mechanisms; and confidence-aware objectives (Burns et al., 2023; Guo et al., 2024) further balance weak guidance with student predictions to avoid overfitting to noisy labels. Inspired by these advancements, we leverage diverse weak teachers to diversify and improve the supervision signals.
3 Weak-to-Strong Learning for VQA
3.1 Problem Setup
Assume that we have access to a weak VQA model $f_{\text{weak}}$, which in practice can be instantiated by existing open-source VQA models. Let $D_{\text{w2s}}=\{x_{1},x_{2},\ldots,x_{n}\}$ denote an unlabeled video dataset with no ground-truth labels. We use $f_{\text{weak}}$ to generate predictions $\hat{y}_{j}=f_{\text{weak}}(x_{j})$ for each video $x_{j}\in D_{\text{w2s}}$, and subsequently train or fine-tune a strong student model $f_{\text{w2s}}$ on $D_{\text{w2s}}$ using these predictions as supervision. The objective is to examine whether $f_{\text{w2s}}$ can outperform $f_{\text{weak}}$ without relying on human annotations for training.
3.2 Weak-to-Strong Implementation for VQA
Weak Models $f_{\text{weak}}$. We select five open-source VQA models [^2] Here, the term “weak” is relative to the student model. In fact, the selected models represent state-of-the-art VQA approaches. $f_{\text{weak}}$: MinimalisticVQA (VII) (Sun et al., 2024), MinimalisticVQA (IX) (Sun et al., 2024), FAST-VQA (Wu et al., 2022), DOVER (Wu et al., 2023a), and Q-Align (Wu et al., 2023b). All models are trained on the LSVQ dataset (Ying et al., 2021) and encompass architectures including convolutional neural networks, vision transformers, and LMMs. Detailed descriptions of these methods are provided in Appendix B.1.
Figure 2: Overview of our weak-to-strong training pipeline.
Strong Model $f_{\text{w2s}}$. For the strong student model, we adopt LLaVA-OneVision-Chat-7B (Li et al., 2024), a LMM whose capacity substantially exceeds that of the weak teachers, as the backbone. A detailed comparison of model parameters and architecture is provided in Table 4. To better adapt it to the VQA task, we follow a preprocessing strategy similar to LMM-VQA (Ge et al., 2025): one key frame per second is sampled for the vision encoder, while motion features are extracted for each key frame using all frames within that second via SlowFast (Feichtenhofer et al., 2019). These motion features are then processed by a motion projector and fused with the visual features before being fed into the language model of the LMM. A detailed description of our student model is provided in Appendix C.1, and its overall architecture is illustrated in Figure 3.
Training Dataset $D_{\text{w2s}}$. We first collect a pool of $3$ million videos from popular social media platforms, including YouTube, TikTok, Youku, and Bilibili. From this pool, we select a subset using a mixed-integer programming approach (Vonikakis et al., 2017) to match the target distributions of LSVQ—the training set of the teacher models—across nine low-level metrics that quantify visual characteristics: blockiness (Romaniak et al., 2012), blur (Narvekar & Karam, 2011), contrast (Peli, 1990), noise, flickering (Pandel, 2008), colorfulness (Hasler & Suesstrunk, 2003), luminance, temporal information, and spatial information (ITU-T P.910, 2008). We then sample 200200k videos from the matched subset to construct a representative and diverse training set for the student model, covering a wide range of quality conditions. A detailed description of the dataset construction procedure and analysis is provided in Appendix A.
Training Protocol. We train $f_{\text{w2s}}$ on $D_{\text{w2s}}$, where supervision is provided by pseudo-labels generated from $f_{\text{weak}}$, and optimize the model with the standard cross-entropy loss. Training is conducted with AdamW, an initial learning rate of $1\times 10^{-4}$, a cosine decay schedule, and a weight decay of $0.05$. We use a batch size of $16$ and train for 200200k iterations with linear warm-up in the first 66k steps. All experiments are implemented in PyTorch and trained on $8$ NVIDIA A800 GPUs over approximately two days.
Validation Datasets. To comprehensively assess model performance, we evaluate on ten VQA benchmarks grouped into in-domain and out-of-distribution (OOD) categories. The in-domain datasets include LSVQ Test (Ying et al., 2021), LSVQ 1080p (Ying et al., 2021), KoNViD-1k (Hosu et al., 2017), LIVE-VQC (Sinno & Bovik, 2018), and YouTube-UGC (Wang et al., 2019), all consisting of user-generated content (UGC) videos. The OOD datasets comprise LIVE-YT-Gaming (Yu et al., 2022), CGVDS (Saha et al., 2023), LIVE-YT-HFR (Madhusudana et al., 2021), Waterloo-IVC-4K (Li et al., 2019b), and KVQ (Lu et al., 2024), which differ from in-domain benchmarks in both content distribution and distortion types. Further details of these datasets are provided in Appendix A.4.
Figure 3: Overall architecture of our strong student model. Following LMM-VQA (Ge et al., 2025), we use a dual-branch visual encoder with an additional motion module for temporal distortion modeling. The model supports both single- and dual-video input strategies with distinct training and inference designs.
Evaluation Metrics. We adopt two widely used criteria to evaluate the performance of VQA models: Spearman Rank Correlation (SRCC) and Pearson Linear Correlation (PLCC), which indicate the prediction monotonicity and prediction linearity, respectively.
3.3 Experimental Results and Analysis
We report overall in-domain and OOD performance in Figure 4, with per-dataset results provided in Appendix D.1. For in-domain benchmarks, the student model achieves performance comparable to its teachers, with an average improvement of $0.24\%$, indicating that our simple knowledge distillation approach effectively transfers quality assessment knowledge from weak to strong models. While for OOD benchmarks, the student exhibits substantial average gains of $7.87\%$ over its teachers, highlighting a pronounced weak-to-strong generalization effect. Interestingly, for stronger teacher models such as MinimalisticVQA (IX) and Q-Align, we observe that their student counterparts achieve comparable performance on in-domain benchmarks and even surpass the supervised models on OOD benchmarks. We attribute this to the larger training dataset (200200k videos), which, although pseudo-labeled by VQA models, elicits stronger generalization capabilities than the human-labeled LSVQ dataset (2727k videos).
In summary, our results empirically demonstrate a clear weak-to-strong generalization effect in VQA, where the most significant improvements arise on OOD data unseen during training. This finding is particularly important for VQA, as in-domain performance on existing benchmarks has largely saturated and even risks overfitting, while current methods suffer from severe degradation on OOD scenarios. Weak-to-strong generalization therefore offers a promising paradigm for addressing this challenge, and in the next section we present a practical solution.
Figure 4: Student model performance under pseudo-labels from five weak models: MinimalisticVQA (VII), MinimalisticVQA (IX), FAST-VQA, DOVER, and Q-Align (left to right).
4 Improving Weak-to-Strong Learning for VQA
We enhance weak-to-strong generalization in VQA from two aspects: (1) unifying diverse supervision signals and (2) iterative W2S training, both aimed at expanding the generalization capacity of the student model.
4.1 Unifying Diverse Supervision Signals
4.1.1 Ranking-based VQA Method
Absolute quality scores obtained from different labeling manners may be inconsistent in their ranges and scales, making them unsuitable for regression-based training. In contrast, the relative quality ranks of video pairs within the same manner are consistent. To unify these heterogeneous supervision signals, we reformulate quality prediction as a ranking problem, enabling the model to learn quality assessment capability through pairwise comparisons.
Specifically, given a video pair $(\bm{x}^{A},\bm{x}^{B})$, we input them into the student model defined in Section 3.2, which is trained to predict their relative quality. Following (Zhu et al., 2024), we adopt ranking labels {“superior”, “better”, “similar”, “worse”, “inferior”} to refine ranking accuracy. During inference, we employ the adaptive soft comparison method (Zhu et al., 2024) to derive quality scores. It first computes a soft probability matrix over ranking categories by comparing each test video against anchor videos, and then applies maximum a posteriori (MAP) estimation (Tsukida et al., 2011) under Thurstone’s Case V model (Thurstone, 2017) to obtain calibrated quality scores. The detailed inference procedure is provided in Appendix C.3.
Figure 5: Our pairwise quality annotations consist of two types: (1) pseudo-labeling based on ensembling homogeneous teachers, and (2) quality ranking derived from integrating heterogeneous teachers.
4.1.2 Ensembling Homogeneous Teachers
In Section 3.3, we observe that stronger teacher models generally yield more capable students, in some cases even surpassing fully supervised counterparts. A naïve strategy is thus to enhance the accuracy of teacher models. To this end, we adopt a simple approach: averaging ensemble predictions from five VQA methods in Section 3.2 to improve the reliability of the supervision signals.

### v4

# Breaking Annotation Barriers: Generalized Video Quality Assessment via Ranking-based Self-Supervision

@@ Abstract
Video quality assessment (VQA) is essential for quantifying perceptual quality in various video processing workflows, spanning from camera capture systems to over-the-top streaming platforms. While recent supervised VQA models have made substantial progress, the reliance on manually annotated datasets—a process that is labor-intensive, costly, and difficult to scale up—has hindered further optimization of their generalization to unseen video content and distortions. To bridge this gap, we introduce a self-supervised learning framework for VQA to learn quality assessment capabilities from large-scale, unlabeled web videos. Our approach leverages a learning-to-rank paradigm to train a large multimodal model (LMM) on video pairs automatically labeled via two manners, including quality pseudo-labeling by existing VQA models and relative quality ranking based on synthetic distortion simulations. Furthermore, we introduce a novel iterative self-improvement training strategy, where the trained model acts an improved annotator to iteratively refine the annotation quality of training data. By training on a dataset $10\times$ larger than the existing VQA benchmarks, our model: (1) achieves zero-shot performance on in-domain VQA benchmarks that matches or surpasses supervised models; (2) demonstrates superior out-of-distribution (OOD) generalization across diverse video content and distortions; and (3) sets a new state-of-the-art when fine-tuned on human-labeled datasets. Extensive experimental results validate the effectiveness of our self-supervised approach in training generalized VQA models. The datasets and code will be publicly released to facilitate future research. Our code and database will be available at https://github.com/clh124/LMM-PVQA.
Figure 1: Overview of our work. (a) Existing state-of-the-art VQA models exhibit poor out-of-distribution performance. (b) We construct a large-scale VQA dataset consisting of 700700k video pairs, sampled from multiple social media platforms, covering more than 20 content categories. (c) We explore two strategies for automatically annotating the relative quality of video pairs. (d) Our proposed model undergoes iterative training to enhance generalization performance.
footnotetext: *Equal contribution. Project lead.
@@ 1 Introduction
Video quality assessment (VQA) [35] This work focuses on no-reference (NR) or blind VQA, which assesses video quality without relying on additional reference information. plays an important role in modern video processing systems, delivering objective quality measurements used to optimize end-user Quality of Experience (QoE). With the advances in deep neural networks (DNNs) [15, 10, 29] and the increasing availability of human-annotated VQA datasets [16, 50, 60, 70], current VQA models [52, 61, 64, 65] have achieved significant progress through supervised learning. Nevertheless, supervised learning inherently faces a limitation: the generalization of the VQA models heavily depends on the diversity of the training data. For example, even top-tier VQA models [52, 61, 64, 65] exhibit significant performance drops in out-of-distribution evaluations, as illustrated in Fig. 1(a).
@@
Therefore, the development of specialized VQA datasets remains critical to address the growing diversity of emerging media formats and their associated distortions [76, 75, 49, 67, 26, 33, 71]. However, constructing such datasets is highly resource-intensive. A standardized subjective experiment comprises two key phases: test sample curation and subjective quality annotation. The test sample curation phase necessitates rigorous selection of representative video samples, as inadequate sampling strategies risk producing oversimplified datasets (i.e.,easy dataset” problem [52, 4]) and may induce model overfitting. Meanwhile, subjective annotation—though vital—is laborious and costly. International Telecommunication Union (ITU) standards (e.g., ITU-T P.910 [18]) outline specific recommendations for experimental setups, including display conditions, stimulus duration, subject count, and rating methodologies. These constraints, though necessary for statistically meaningful results, impede large-scale dataset expansion due to prohibitive annotation costs.
@@
Self-supervised or unsupervised learning [13], which eliminates the need for human-labeled annotations, is a potential solution in VQA to mitigate the high costs of subjective experiments. However, current self-supervised VQA methods [7, 6, 8, 34, 36] still lag behind their supervised counterparts in performance. Typical implementations of current self-supervised VQA methods [6, 34] employ contrastive learning frameworks with proxy tasks such as distortion type/severity classification on synthetically generated data. However, it suffers from two shortcomings: (1) they fail to capture visual content and aesthetic characteristics relevant to perceptual quality assessment, and (2) they inadequately model real-world authentic distortion patterns that follow complex nonlinear degradation processes.
In this paper, we propose a novel self-supervised learning framework for VQA to address the aforementioned challenges. Specifically, we reformulate quality regression as a ranking problem, enabling the model to learn quality assessment capabilities through pairwise comparisons. This approach is motivated by the observation that pairwise ranking is more intuitive and reliable than absolute quality ratings for human evaluators. We then explore two strategies to automatically label the relative quality of video pairs. The first leverages existing VQA model as the “judge” to determine the relative quality, where we ensemble the results from multiple judges to mitigate inherent evaluation biases in individual models. This process is iterative: once a new VQA model is trained, it can serve as an enhanced judge, achieving self-improvement through repeated refinement. The second approach relies on synthetic distortion simulations, where we introduce various types of distortions at different severity levels and utilize these severity levels to establish relative quality rankings.
We adopt a large multimodal model (LMM) integrated with a fixed motion encoder and a learnable motion projector as our VQA model. Given two videos and a text prompt as input, the model outputs a textual judgment indicating whether the first video has higher or lower quality than the second. We employ a standard supervised fine-tuning (SFT) strategy to train the proposed VQA model using automatically annotated video pairs. Training is conducted in three iterative stages with progressively increasing sample sizes of 500500k, 600600k, and 700700k video pairs, respectively. We evaluate the proposed model on ten diverse VQA benchmarks. Experimental results show that the model achieves zero-shot performance comparable to, or even surpassing, existing supervised VQA models, while demonstrating superior out-of-distribution generalization. Moreover, after fine-tuning on human-annotated datasets, the model significantly outperforms state-of-the-art supervised approaches.
@@
  • We construct a large-scale VQA dataset comprising 700k video pairs with both authentic and synthetic distortions, each annotated with automatically generated quality labels.
  • We propose a self-supervised VQA framework that enables the model to learn quality assessment capabilities through pairwise comparisons, enhanced by an iterative training strategy for continuous self-improvement.
  • Our model achieves strong zero-shot performance across multiple VQA benchmarks, highlighting its effectiveness and generalization in video quality assessment.
@@ 2.1 VQA Models
Supervised VQA. Early supervised VQA models, such as V-BLIINDS [46], TLVQM [22], and VIDEVAL [55], are typically knowledge-driven. These methods extract quality-related features based on natural scene statistics (NSS) [37], blurriness [22], motion vectors [21], optical flow [3], and other handcrafted cues, followed by training a machine learning-based regressor, such as a support vector regressor (SVR) or random forest regressor, to derive quality scores.
With the popularity of DNNs, some VQA methods, such as VSFA [25], Li22 [23], PatchVQA [70], etc., employ pre-trained DNNs as feature extractors to derive quality representations from video frames, followed by training quality regressors to map the extracted features into quality scores. Commonly used feature extractors include image classification models [15], image quality assessment (IQA) models [73, 69], and action recognition models [12], while commonly used quality regressors often consist of GRUs [25], Transformers [62], and InceptionTime [70], etc.
Many recent works explore the use of 2D or 3D convolutional neural networks (CNNs) [15], Vision Transformers (ViTs) [10], and LMMs [65] as feature extractors, fine-tuning them in an end-to-end manner to achieve state-of-the-art performance. For instance, SimpleVQA [51] and MinimalisticVQA [52] perform end-to-end training of the spatial feature extractor (e.g., ResNet [15], Swin [29]) while adopting a temporal extractor (i.e., SlowFast [12]) with fixed pre-trained weights. FAST-VQA [61] trains a 3D DNN (i.e., VideoSwin [31]) in an end-to-end fashion, and DOVER [64] further extends FAST-VQA with an aesthetic evaluation branch based on a 2D CNN (i.e., ConvNet [30]). Q-Align [65] fine-tunes the visual encoder of an LMM with a predefined text-based quality inquiry prompt.
Unsupervised and self-supervised VQA. A class of popular unsupervised or self-supervised VQA approaches [6, 34, 36, 66] aims to learn quality-aware feature representations from scratch, followed by fine-tuning a linear projector with human-annotated labels to serve as a quality score regressor. For example, CSPT [6] and CONVIQT [34] adopt contrastive learning frameworks with proxy tasks such as next-frame feature discrimination and distortion type/severity classification to learn the quality-aware representation. QPT V2 [66], on the other hand, employs an encoder-decoder architecture to reconstruct pristine videos from distorted ones, thereby learning distortion-aware representations.
Another line of research focuses on developing opinion-unaware VQA methods, which aligns with the objective of this work. Some knowledge-driven VQA methods [38, 72, 27, 39, 20] estimate video quality by directly measuring the distribution distance of specific features between distorted and pristine videos. For instance, NIQE [38] assesses the spatial quality of a test image by computing the distance between the Multivariate Gaussian model fitted to local features of the test image and pristine natural images. TPQI [27] captures temporal distortions by analyzing the straightness and compactness of video trajectories within perceptual domains of the human visual system. Recent studies [64, 1] leverage visual-language models to achieve zero-shot I/VQA. For example, BUONA-VISTA [63] employs CLIP [44] to estimate the relatively probabilities of text promt “high quality” compared to “low quality”, and combines these with NIQE and TPQI scores to assess video quality.
2.2 VQA Datasets
The progress of data-driven VQA models heavily relies on VQA datasets annotated by human subjects. Early VQA datasets, such as LIVE-VQA [48] and CSIQ-VQA [57], primarily focus on compression and transmission distortions in professionally generated content (PGC). However, these datasets contain a limited number of source videos and only a few hundred distorted samples. While they have contributed to the development of knowledge-driven VQA methods, their scale and diversity make them inadequate for training data-driven VQA models. With the rise of social media platforms, the focus has shifted to user generated content (UGC) videos characterized by natural and in-the-wild distortions. Current mainstream UGC VQA datasets, such as KoNVid-1k [16], YouTube-UGC [60], and LSVQ [70], contain thousands to tens of thousands of videos, providing large-scale benchmarks that have significantly advanced research in data-driven VQA. Meanwhile, with the emergence of new media formats, specialized VQA datasets have been developed to address domain-specific challenges, such as gaming videos [71], 360-degree videos [67], 4K videos  [26], high-frame-rate videos [33], AIGC videos [75], etc., each designed to facilitate the development of expert VQA models tailored to specific video types.
3 Pairwise-Labeled Video Dataset
This section describes the construction of the large-scale pairwise-labeled video dataset (PLVD) in detail.
Figure 2: Distribution of nine metrics across mainstream UGC datasets (LSVQ, KoNVid-1k, YouTube-UGC, LIVE-VQC), as well as our dataset before and after sampling.
3.1 Video Collection
Source Video Collection. We create a large-scale dataset comprising $3$ million videos collected from popular social media platforms, including YouTube, TikTok, Youku, and Bilibili. This dataset encompasses a wide range of content categories and distortion scenarios, such as vlogs, gaming videos, animations, live streams, etc., ensuring a representative and diverse collection with varying quality conditions. We provide the detail analysis of source videos in the supplement file.
Candidate Video Sampling. We select a subset of videos from the collected pool using a mixed-integer programming approach [56] to match target distributions defined by nine low-level metrics that quantify visual characteristics of datasets, including blockiness [45], blur [41], contrast [43], noise, flickering [42], colorfulness [14], luminance, temporal information (TI) [19], and spatial information (SI) [19]. We provide the calculation of these metrics in the supplement file. Our target distribution mirrors the aggregated distributions of mainstreaming UGC datasets—LSVQ [70], KonViD-1k [16], YouTube-UGC [60], and LIVE-VQC [50]—to ensure compatibility with existing benchmarks. Finally, we sample 438438k videos to enhance diversity in content and scenes while maintaining close alignment with mainstream UGC datasets. As illustrated in Fig. 2, the distributions of the sampled subset exhibit strong consistency with that of public VQA datasets across all nine metrics.
3.2 Pairwise Quality Annotation
We explore two strategies to automatically annotate the relative quality of video pairs: (1) quality pseudo-labeling using existing VQA models, and (2) relative quality ranking based on synthetic distortion simulations. We define two types of ranking labels: hard ranking and soft ranking. Specifically, hard ranking consists of two categories, “better” and “worse”, while soft ranking provides finer-grained comparisons with five levels: “superior”, “better”, “similar”, “worse”, and “inferior”, to reflect varying degrees of relative quality.
Table 1: Statistics of raw videos and video pairs in the PLVD dataset. The values in each cell indicate the number of videos or video pairs in PLVD-Part1/-Part2/-Part3
Videos Video Pairs
PLVD-VQA 200k/100k/50k 250k/85k/85k
PLVD-DS Spatial 50k/2k/2k 160k/5k/5k
Temporal 20k/1k/1k 40k/5k/5k
Compression 10k/1k/1k 50k/5k/5k
Total 280k/384k/438k 500k/600k/700k
3.2.1 Pseudo-labeling using VQA Models
Inspired by recent LLM-as-a-Judge methods [58] that leverage large language models (LLM) to model user preference for LLM alignment, we propose utilizing established VQA models as judges to evaluate the relative quality of video pairs. We refer to this subset of data as PLVD-VQA. Specifically, we choose five state-of-the-art VQA models: Minimalistic-VQA (VII) [52], Minimalistic-VQA (IX) [52], FAST-VQA [61], DOVER [64], and Q-Align [65], all trained on LSVQ [70], as our initial judges. For a video pair $(x^{A},x^{B})$, each judge model $\mathcal{J}_{i}$ generates quality scores $j^{A}_{i}$ and $j^{B}_{i}$. For hard ranking, the judgment is determined by comparing the mean quality scores $\overline{j}^{A}$ and $\overline{j}^{B}$. If $\overline{j}^{A}>\overline{j}^{B}$, we annotate $x^{A}$ as higher quality than $x^{B}$ (label: “better”); otherwise, $x^{A}$ is labeled as worse quality (label: “worse”).